# Model-free policy gradient for discrete-time mean-field control

### v1

@@ 2.3 Perturbed Process and Value Function
We first give an informal presentation of the perturbation argument. For $\pi\in\Pi,\varepsilon>0$ and a sequence of perturbations $\lambda$ == $(\lambda_{t})_{t=0}^{T-1}$, we consider processes $X^{\pi}_{t},Y^{\pi,\lambda}_{t}$ such that
@@
where $l_{t}^{\pi}\coloneqq\operatorname*{logit}(\mathbb{P}_{X_{t}^{\pi}})$. We consider an i.i.d. sequence of random perturbations $\Lambda=(\Lambda_{t})_{t=0}^{T-1}$ of the state distribution at each time step, independent of the idiosyncratic noise, and distributed according to $\nu$. As $\mathcal{P}(\mathcal{X})$ is not a vector space, one needs to be careful about the definition of $(\Lambda_{t})$. Thanks to the logits parametrization introduced in Section 2.1, this perturbation can easily be applied in the space of logits. We write $\tilde{\pi}({\mathrm{d}}a\mid t,x,l)=\pi({\mathrm{d}}a\mid t,x,\operatorname*{softmax}(l))$. Then, for a fixed $\pi$, $Y^{\pi,\Lambda}$ is a process for which the distributional dependences are randomized by $\Lambda$. We consider the following approximation of the value function for a policy $\pi$
$V_{\varepsilon}^{\pi}(\mu_{0})\coloneqq\mathbb{E}^{\pi}_{\xi\sim\mu_{0}}\left[\sum_{t=0}^{T-1}r(Y_{t}^{\pi,\Lambda},\alpha^{\pi,\Lambda}_{t},\operatorname*{softmax}(l_{t}^{\pi}+\varepsilon\Lambda_{t}))+g(Y^{\pi,\Lambda}_{T},\operatorname*{softmax}(l_{T}^{\pi}+\varepsilon\Lambda_{T}))\right].$
@@
$V_{\varepsilon}^{\pi}(\mu_{0})$ $=\int\left(\sum_{t=0}^{T-1}\tilde{r}(y_{t},a_{t},l^{\pi}_{t}+\varepsilon\lambda_{t})+\tilde{g}(y_{T},l_{T}^{\pi}+\varepsilon\lambda_{T})\right)\mu_{0}({\mathrm{d}}y_{0})\cdot$
$\qquad\prod_{\tau=0}^{T-1}\nu({\mathrm{d}}\lambda_{\tau})\tilde{\pi}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},l^{\pi}_{\tau}+\varepsilon\lambda_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},l^{\pi}_{\tau}+\varepsilon\lambda_{\tau}).$
@@ 2.4 Derivation and Convergence of Perturbed Gradient
We now present a policy gradient formula for the perturbed value function $V_{\varepsilon}(\mu,\theta)=V_{\varepsilon}^{\pi_{\theta}}(\mu)$. It is the foundation of our model-free estimator of $\nabla_{\theta}V(\mu,\theta)$.
@@
Let Assumption 1 hold. Let $\Lambda_{t},t=0,\ldots,T-1$ be i.i.d. $\mathcal{N}(0,I_{d})$ random variables. Assume that for any $\pi_{\theta}$, $(X_{t}^{\theta})_{t=0}^{T}$ is an $\mathcal{X}$-valued process independent of $(\Lambda_{t})_{t=0}^{T-1}$ and $(Y_{t}^{\theta,\varepsilon})_{t=0}^{T}$ is an $\mathcal{X}$-valued process such that, conditioned on $\Lambda=\lambda=(\lambda_{0},\ldots,\lambda_{T-1})\in{\mathbb{R}}^{dT}$, $X_{t}^{\theta}$ and $Y^{\theta,\varepsilon}$ are distributed according to (2.3) with $\pi=\pi_{\theta}$ and initial condition $\xi$ such that $\mu_{0}=\mathcal{L}(\xi)\in\mathcal{P}(\mathcal{X})^{*}$. Let $(\alpha_{t}^{\theta})_{t=0}^{T-1}$ and $(\alpha_{t}^{\theta,\varepsilon})_{t=0}^{T-1}$ be the corresponding controls. Then, for $\theta\in\Theta$, the perturbed value function $V_{\varepsilon}(\mu_{0},\theta)=V_{\varepsilon}^{\pi_{\theta}}(\mu_{0})$ is differentiable in $\theta$ with
$\nabla_{\theta}V_{\varepsilon}(\mu_{0},\theta)=\mathbb{E}\Bigg[\sum_{t=0}^{T}\left(\varepsilon^{-1}\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}+1_{t<T}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t}\right)G_{t}^{\varepsilon}\Bigg],$ (2.5)
@@
Notice that, contrary to the exact policy gradient formula (2.1), the formula (2.5) only involves derivatives with respect to the trainable parameters $\theta$. However, estimating this quantity numerically is not straightforward due to the gradient of logits $\nabla_{\theta}l_{t}^{\theta}$. To overcome this challenge, one can use the same perturbation method, now applied to a problem with the same dynamics but with a different reward, as detailed in Section 3.1,
@@
We want to highlight that both the gradient representation of Theorem 2.3 and the convergence result Theorem 2.4 hold independently of the way the processes $X^{\theta}$ and $Y^{\theta,\Lambda}$ are generated, as long as the assumptions of the theorems are satisfied. In particular, we do not make any coupling assumption between $X_{t}^{\theta}$ and $Y_{t}^{\theta,\Lambda}$. In practice, using coupled samples might help reduce variance in the gradient estimates when subtracting a baseline depending on $X_{t}^{\theta}$ to the cumulated rewards in (2.5), as commonly done in single-agent REINFORCE implementations [44, 39].
@@ 3.1 Model-Free Estimator
Notice that we used the causality and log-likelihood trick: we are able to remove the terms $s\geq t$ due to the ‘reward’ at time $t$ not depending on the actions at time $s\geq t$. Indeed, for $s\geq t$, by denoting $\mathcal{F}_{s}$ the filtration generated by $(Y_{\tau\leq s}^{\theta,\varepsilon},\Lambda_{\tau\leq s},\alpha_{\tau<s}^{\theta,\varepsilon})$, we have
@@
$=\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\right]\mathbb{E}\left[\varepsilon^{-1}\Lambda_{s}\nabla_{\theta}l_{s}^{\theta}\right]+\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\varepsilon\Lambda_{s})\right]$
$=\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\right]\cdot 0+\mathbb{E}\left[\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\epsilon\Lambda_{s})\mid\mathcal{F}_{s}\right]\right]$
$=0+\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\int\nabla_{\theta}\tilde{p}(\theta,a,s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\epsilon\Lambda_{s})\,da\right]$
$=\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\nabla_{\theta}\left(\int\tilde{p}(\theta,a,s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\epsilon\Lambda_{s})\,da\right)\right]$
$=\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\nabla_{\theta}(1)\right]=0.$
Hence, assuming we have access to $\mathbb{P}(X_{t}^{\theta}=x^{(i)})$ for every $i=1,\ldots,d$ (either via direct observation or by sampling many particles), we get $T\times d\times D$ linear equations with unknowns $\frac{\partial}{\partial\theta_{k}}\mathbb{P}(X_{t}^{\theta}=x^{(i)})$, for $t=0,\ldots,T-1,i=1,\ldots,d,k=1,\ldots,D$, where the coefficients can be estimated by Monte Carlo sampling. This system is triangular in the time variable, and it is easy to see that $\nabla_{\theta}\mathbb{P}(X_{0}^{\theta}=x^{(i)})=0$ for all $i$, hence the system can be solved by forward substitution. Still, this estimator is biased due to the approximation in the perturbation method, but the bias can be made arbitrarily small by taking $\varepsilon$ small enough. The bias of this estimator is studied in Section 3.3. However, taking $\varepsilon$ too small may lead to high variance in the estimator. In practice, the linear system might be easier to solve by subtracting a baseline from the first term
@@ 3.2 Algorithm
7:   Sample $\Lambda_{t}\sim\mathcal{N}(0,I_{d})$ independently for $t=0,\ldots,T-1$.
@@
22:  $\widehat{\nabla V}(\theta)\leftarrow\frac{1}{k}\sum_{k=1}^{n}\widehat{\nabla V}(\theta)^{(k)}$.
@@
Notice that for every time step $t=1,\ldots,T$ we resample new trajectories to $\nabla_{\theta}l_{t}^{\theta}$ in Algorithm 2, in order to reduce the bias induced by the approximation of $\nabla_{\theta}l_{s}^{\theta}$ for $s\leq t-1$. Similarly, we use independent trajectories for estimating $\nabla_{\theta}l_{t}^{\theta}$ and $\nabla_{\theta}V_{\varepsilon}(\mu,\theta)$.
@@ 3.3 Bias of Estimator
Moreover, $\tilde{P}$ and $\tilde{p}$ are twice differentiable in $l\in{\mathbb{R}}^{d}$, and there exists $B_{1},B_{2}\geq 0$ such that for all $\theta\in\Theta,x,x^{\prime}\in\mathcal{X},a\in\mathcal{A}$ and $l\in{\mathbb{R}}^{d}$,
$\left|\frac{\partial}{\partial l_{i}}\left(\tilde{P}(x^{\prime}\mid x,a,l)\tilde{p}(\theta,a,x,l)\right)\right|\leq B_{1},\>\forall i,$
$\left|\frac{\partial^{2}}{\partial l_{i}\partial l_{j}}\left(\tilde{P}(x^{\prime}\mid x,a,l)\tilde{p}(\theta,a,x,l)\right)\right|\leq B_{2},\>\forall i,j.$
@@
$\beta\coloneqq\frac{1}{2}\left(1+B_{1}|\mathcal{A}|+\sqrt{(1+B_{1}|\mathcal{A}|)^{2}+4B_{1}|\mathcal{A}|(d-1)}\right).$ (3.3)
@@
Let $\theta\in\Theta$ and $\varepsilon\in(0,1)$. Let $(\Lambda_{t})_{t=0}^{T}$ be i.i.d. $\mathcal{N}(0,I_{d})$ random variables. Let $(X_{t}^{\theta})_{t=0}^{T}$ and $(Y_{t}^{\theta,\varepsilon})_{t=0}^{T}$ be $\mathcal{X}$-valued processes such that, conditionally on $\Lambda=(\Lambda_{0},\Lambda_{1},\ldots,\Lambda_{T})$, they are distributed according to the perturbed dynamics (2.3) with policy $\pi=\pi_{\theta}$ and initial condition $\xi$ satisfying $\mu_{0}=\mathcal{L}(\xi)\in\mathcal{P}(\mathcal{X})^{*}$. Let $(\alpha_{t}^{\theta})_{t=0}^{T-1}$ and $(\alpha_{t}^{\theta,\varepsilon})_{t=0}^{T-1}$ denote the associated controls of $X^{\theta}$ and $Y^{\theta,\varepsilon}$.
@@
Let Assumptions 1 and 2 hold. Let $n\in{\mathbb{N}}$ and $\varepsilon\in(0,1)$. Under the sampling and perturbation setup above, let $\{(X^{\theta,t,k},Y^{\theta,\varepsilon,t,k},\Lambda^{t,k},\alpha^{\theta,t,k},\alpha^{\theta,\varepsilon,t,k})\}_{k=1}^{n}$ be i.i.d. copies truncated at time $t$, for all $t=1,\ldots,T$. Moreover, assume that the trajectories $(X^{\theta,t,k}_{\leq t},Y^{\theta,\varepsilon,t,k}_{\leq t},\Lambda^{t,k}_{<t},)$ for $k=1,\ldots,n$ and $t=1,\ldots,T$ are independent. Define $\widehat{\nabla_{\theta}l^{\theta}_{0}}=0\in{\mathbb{R}}^{d\times D}$ and, for $t=1,\ldots,T$,
$\widehat{\nabla_{\theta}\mu_{t}^{(i)}}$ $=\frac{1}{n}\sum_{k=1}^{n}1_{Y^{\theta,\varepsilon,t,k}_{t}=x^{(i)}}\left(\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}^{t,k}\widehat{\nabla_{\theta}l^{\theta}_{s}}+\nabla_{\theta}\log\tilde{\pi}_{\theta}\left(\alpha_{s}^{\theta,\varepsilon,t,k}\mid Y_{s}^{\theta,\varepsilon,t,k},l_{s}^{\theta}+\varepsilon\Lambda_{s}^{t,k}\right)\right),$ (3.4)
@@
$E_{t}(\theta,\varepsilon)\coloneqq\left\|\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]-\nabla_{\theta}l_{t}^{\theta}\right\|\leq\mathbf{\varepsilon}\left((K_{1,\varepsilon}+K_{2,\varepsilon})\beta_{\varepsilon}^{t}-K_{1,\varepsilon}\beta^{t}-K_{2,\varepsilon}\right),$ (3.5)
where $\beta_{\varepsilon},K_{1,\varepsilon},K_{2,\varepsilon}$ are constants defined in (5.13) depending on $\varepsilon$ and the problem parameters, bounded as $\varepsilon\to 0$. Otherwise, $\left\|\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]-\nabla_{\theta}l_{t}^{\theta}\right\|\leq\mathbf{s}^{\theta}\frac{L_{p}^{\prime}}{2}\varepsilon t$ where $\mathbf{s}^{\theta}\coloneqq\sup_{t\geq 0}\sum_{i=1}^{d}\mathbb{P}(X_{t}^{\theta}=x^{(i)})^{-1}$.
@@
The reward functions $r$ and $g$ are globally Lipschitz in the measure argument, i.e., there exists constants $L_{r},L_{g}>0$ such that, for all $x\in\mathcal{X},a\in\mathcal{A},\mu,\mu^{\prime}\in\mathcal{P}(\mathcal{X})$,
@@
Moreover, $\tilde{r}$ and $\tilde{g}$ are twice differentiable in $l\in{\mathbb{R}}^{d}$, and there exists $M_{1},M_{2}>0$ such that for all $x\in\mathcal{X},a\in\mathcal{A}$ and $l\in{\mathbb{R}}^{d}$,

### v2

@@ 2.3 Perturbed Process and Value Function
We first give an informal presentation of the perturbation argument. For $\pi\in\Pi,\varepsilon>0,$ and a sequence of perturbations $\lambda$ == $(\lambda_{t})_{t=0}^{T}$, we consider processes $X^{\pi}_{t},Y^{\pi,\lambda}_{t}$ such that
@@
where $l_{t}^{\pi}\coloneq\operatorname*{logit}(\mathbb{P}_{X_{t}^{\pi}})$. We consider an i.i.d. sequence of random perturbations $\Lambda=(\Lambda_{t})_{t=0}^{T}$ of the state distribution at each time step, independent of the idiosyncratic noise, and distributed according to $\nu$. As $\mathcal{P}(\mathcal{X})$ is not a vector space, one needs to be careful about the definition of $(\Lambda_{t})$. Thanks to the logits parametrization introduced in Section 2.1, this perturbation can easily be applied in the space of logits. We write $\tilde{\pi}({\mathrm{d}}a\mid t,x,l)=\pi({\mathrm{d}}a\mid t,x,\operatorname*{softmax}(l))$. Then, for a fixed $\pi$, $Y^{\pi,\Lambda}$ is a process for which the distributional dependences are randomized by $\Lambda$. We consider the following approximation of the value function for a policy $\pi$
$V_{\varepsilon}^{\pi}(\mu_{0})\coloneq\mathbb{E}^{\pi}_{\xi\sim\mu_{0}}\left[\sum_{t=0}^{T-1}r(Y_{t}^{\pi,\Lambda},\alpha^{\pi,\Lambda}_{t},\operatorname*{softmax}(l_{t}^{\pi}+\varepsilon\Lambda_{t}))+g(Y^{\pi,\Lambda}_{T},\operatorname*{softmax}(l_{T}^{\pi}+\varepsilon\Lambda_{T}))\right].$ (2.4)
@@
$V_{\varepsilon}^{\pi}(\mu_{0})$ $=\int\left(\sum_{t=0}^{T-1}\tilde{r}(y_{t},a_{t},l^{\pi}_{t}+\varepsilon\lambda_{t})+\tilde{g}(y_{T},l_{T}^{\pi}+\varepsilon\lambda_{T})\right)\mu_{0}({\mathrm{d}}y_{0})$
$\qquad\cdot\prod_{\tau=0}^{T-1}\nu({\mathrm{d}}\lambda_{\tau})\tilde{\pi}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},l^{\pi}_{\tau}+\varepsilon\lambda_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},l^{\pi}_{\tau}+\varepsilon\lambda_{\tau}).$
@@ 2.4 Derivation and Convergence of Perturbed Gradient
We now present a policy gradient formula for the perturbed value function $V_{\varepsilon}(\mu,\theta)=V_{\varepsilon}^{\pi_{\theta}}(\mu)$ in (2.4). It is the foundation of our model-free estimator of $\nabla_{\theta}V(\mu,\theta)$.
@@
Let Assumption 1 hold. Let $\Lambda_{t},t=0,\ldots,T$ be i.i.d. $\mathcal{N}(0,I_{d})$ random variables. Assume that for any $\pi_{\theta}$, $(X_{t}^{\theta})_{t=0}^{T}$ is an $\mathcal{X}$-valued process independent of $(\Lambda_{t})_{t=0}^{T}$ and $(Y_{t}^{\theta,\varepsilon})_{t=0}^{T}$ is an $\mathcal{X}$-valued process such that, conditioned on $\Lambda=\lambda=(\lambda_{0},\ldots,\lambda_{T})\in{\mathbb{R}}^{dT}$, $X^{\theta}$ and $Y^{\theta,\varepsilon}$ are distributed according to (2.3) with $\pi=\pi_{\theta}$ and initial condition $\xi$ such that $\mu_{0}=\mathcal{L}(\xi)\in\mathcal{P}(\mathcal{X})^{*}$. Let $(\alpha_{t}^{\theta})_{t=0}^{T-1}$ and $(\alpha_{t}^{\theta,\varepsilon})_{t=0}^{T-1}$ be the corresponding controls. Then, for $\theta\in\Theta$, the perturbed value function $V_{\varepsilon}(\mu_{0},\theta)=V_{\varepsilon}^{\pi_{\theta}}(\mu_{0})$ is differentiable in $\theta$ with
$\nabla_{\theta}V_{\varepsilon}(\mu_{0},\theta)=\mathbb{E}\Bigg[\sum_{t=0}^{T}\left(\varepsilon^{-1}\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}+1_{t<T}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t})\right)G_{t}^{\varepsilon}\Bigg],$ (2.6)
@@
Notice that, contrary to the exact policy gradient formula (2.1), the formula (2.6) only involves derivatives with respect to the trainable parameters $\theta$. However, estimating this quantity numerically is not straightforward due to the gradient of logits $\nabla_{\theta}l_{t}^{\theta}$. To overcome this challenge, one can use the same perturbation method applied to a problem with the same dynamics but with a different reward, as detailed in Section 3.1,
@@
We want to highlight that both the gradient representation of Theorem 2.3 and the convergence result Theorem 2.4 hold independently of the way the processes $X^{\theta}$ and $Y^{\theta,\Lambda}$ are generated, as long as the assumptions of the theorems are satisfied. In particular, we do not make any coupling assumption between $X_{t}^{\theta}$ and $Y_{t}^{\theta,\Lambda}$. In practice, using coupled samples might help reduce variance in the gradient estimates when subtracting a baseline depending on $X_{t}^{\theta}$ to the cumulated rewards in (2.6), as commonly done in single-agent REINFORCE implementations ([44, 39]).
@@ 3.1 Model-Free Estimator
Notice that we used the causality and log-likelihood trick: we are able to remove the terms $s\geq t$ due to the ‘reward’ at time $t$ not depending on the actions at time $s\geq t$. Indeed, for $s\geq t$, by denoting $\mathcal{F}_{s}$ the filtration generated by $(Y_{\leq s}^{\theta,\varepsilon},\Lambda_{\leq s},\alpha_{<s}^{\theta,\varepsilon})$, we have
@@
$\qquad=\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\right]\mathbb{E}\left[\varepsilon^{-1}\Lambda_{s}\nabla_{\theta}l_{s}^{\theta}\right]+\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\varepsilon\Lambda_{s})\right]$
$\qquad=\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\right]\cdot 0+\mathbb{E}\left[\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\epsilon\Lambda_{s})\mid\mathcal{F}_{s}\right]\right]$
$\qquad=0+\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\int\nabla_{\theta}\tilde{p}(\theta,a,s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\epsilon\Lambda_{s})\,da\right]$
$\qquad=\mathbb{E}\left[1_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\nabla_{\theta}(1)\right]=0.$
Hence, assuming we have access to $\mathbb{P}(X_{t}^{\theta}=x^{(i)})$ for every $i=1,\ldots,d$ (either via direct observation or by sampling many particles), we get $T\times d\times D$ linear equations with unknowns $\frac{\partial}{\partial\theta_{k}}\mathbb{P}(X_{t}^{\theta}=x^{(i)})$, for $t=0,\ldots,T-1,i=1,\ldots,d,k=1,\ldots,D$, where the coefficients can be estimated by Monte Carlo sampling. This system is triangular in the time variable, and it is easy to see that $\nabla_{\theta}\mathbb{P}(X_{0}^{\theta}=x^{(i)})=0$ for all $i$, hence the system can be solved by forward substitution. Still, the resulting estimator is biased due to the approximation in the perturbation method, but the bias can be made arbitrarily small by taking $\varepsilon$ small enough. However, taking $\varepsilon$ too small may lead to high variance in the estimator. The bias and variance of this estimator are studied in Sections 3.3 and 3.4 respectively. In practice, the linear system might be easier to solve by subtracting a baseline from the first term
@@ 3.2 Algorithm
7:   Sample $\Lambda_{t}\sim\mathcal{N}(0,I_{d})$ independently for $t=0,\ldots,T$.
@@
22:  $\widehat{\nabla V}(\theta)\leftarrow\frac{1}{N}\sum_{k=1}^{N}\widehat{\nabla V}(\theta)^{(k)}$.
@@
Notice that for every time step $t=1,\ldots,T$ we resample new trajectories to estimate $\nabla_{\theta}l_{t}^{\theta}$ in Algorithm 2, in order to reduce the bias induced by the approximation of $\nabla_{\theta}l_{s}^{\theta}$ for $s\leq t-1$. Similarly, we use independent trajectories for estimating $\nabla_{\theta}l_{t}^{\theta}$ and $\nabla_{\theta}V_{\varepsilon}(\mu,\theta)$.
@@ 3.3 Bias of Estimator
Moreover, $\tilde{P}$ and $\tilde{p}$ are twice differentiable in $l\in{\mathbb{R}}^{d}$, and there exist $B_{1},B_{2}\geq 0$ such that for all $\theta\in\Theta,x,x^{\prime}\in\mathcal{X},a\in\mathcal{A},t\in\{0,\ldots,T-1\}$ and $l\in{\mathbb{R}}^{d}$,
$\left|\frac{\partial}{\partial l_{i}}\left(\tilde{P}(x^{\prime}\mid x,a,l)\tilde{p}(\theta,a,t,x,l)\right)\right|\leq B_{1},\>\forall i,$
$\left|\frac{\partial^{2}}{\partial l_{i}\partial l_{j}}\left(\tilde{P}(x^{\prime}\mid x,a,l)\tilde{p}(\theta,a,t,x,l)\right)\right|\leq B_{2},\>\forall i,j.$
@@
$\beta\coloneq\frac{1}{2}\left(1+B_{1}|\mathcal{A}|+\sqrt{(1+B_{1}|\mathcal{A}|)^{2}+4B_{1}|\mathcal{A}|(d-1)}\right).$ (3.3)
@@
Let $\theta\in\Theta$ and $\varepsilon\in(0,1)$. Let $(\Lambda_{t})_{t=0}^{T}$ be i.i.d. $\mathcal{N}(0,I_{d})$ random variables. Let $(X_{t}^{\theta})_{t=0}^{T}$ and $(Y_{t}^{\theta,\varepsilon})_{t=0}^{T}$ be $\mathcal{X}$-valued processes such that, conditionally on $\Lambda=(\Lambda_{0},\Lambda_{1},\ldots,\Lambda_{T})$, they are distributed according to the perturbed dynamics (2.3) with policy $\pi=\pi_{\theta}$ and initial condition $\xi$ satisfying $\mu_{0}=\mathcal{L}(\xi)\in\mathcal{P}(\mathcal{X})^{*}$. Let $(\alpha_{t}^{\theta})_{t=0}^{T-1}$ and $(\alpha_{t}^{\theta,\varepsilon})_{t=0}^{T-1}$ denote the associated controls of $X^{\theta}$ and $Y^{\theta,\varepsilon}$ respectively.
@@
Let Assumptions 1 and 2 hold. Let $n\in{\mathbb{N}}$ and $\varepsilon\in(0,1)$. Under the sampling and perturbation setup above, let $\{(X^{\theta,t,k},Y^{\theta,\varepsilon,t,k},\Lambda^{t,k},\alpha^{\theta,t,k},\alpha^{\theta,\varepsilon,t,k})\}_{k=1}^{n}$ be i.i.d. copies truncated at time $t$, for all $t=1,\ldots,T$. Moreover, assume that the trajectories $(X^{\theta,t,k}_{\leq t},Y^{\theta,\varepsilon,t,k}_{\leq t},\Lambda^{t,k}_{\leq t},)$ for $k=1,\ldots,n$ and $t=1,\ldots,T$ are independent. Define $\widehat{\nabla_{\theta}l^{\theta}_{0}}=0\in{\mathbb{R}}^{d\times D}$ and, for $t=1,\ldots,T$,
$\widehat{\nabla_{\theta}\mu_{t}^{(i)}}$ $=\frac{1}{n}\sum_{k=1}^{n}1_{Y^{\theta,\varepsilon,t,k}_{t}=x^{(i)}}\left(\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}^{t,k}\widehat{\nabla_{\theta}l^{\theta}_{s}}+\nabla_{\theta}\log\tilde{\pi}_{\theta}\left(\alpha_{s}^{\theta,\varepsilon,t,k}\mid s,Y_{s}^{\theta,\varepsilon,t,k},l_{s}^{\theta}+\varepsilon\Lambda_{s}^{t,k}\right)\right),$ (3.4)
@@
$E_{t}(\theta,\varepsilon)\coloneq\left\|\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]-\nabla_{\theta}l_{t}^{\theta}\right\|\leq\mathbf{\varepsilon}\left((K_{1,\varepsilon}+K_{2,\varepsilon})\beta_{\varepsilon}^{t}-K_{1,\varepsilon}\beta^{t}-K_{2,\varepsilon}\right),$ (3.5)
where $\beta_{\varepsilon},K_{1,\varepsilon},K_{2,\varepsilon}$ are constants defined in (5.13) depending on $\varepsilon$ and the problem parameters, bounded as $\varepsilon\to 0$. Otherwise, $\left\|\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]-\nabla_{\theta}l_{t}^{\theta}\right\|\leq\mathbf{s}^{\theta}\frac{L_{p}^{\prime}}{2}\varepsilon t$ where $\mathbf{s}^{\theta}\coloneq\sup_{t\geq 0}\sum_{i=1}^{d}\mathbb{P}(X_{t}^{\theta}=x^{(i)})^{-1}$.
@@
The reward functions $r$ and $g$ are globally Lipschitz in the measure argument, i.e., there exist constants $L_{r},L_{g}>0$ such that, for all $x\in\mathcal{X},a\in\mathcal{A},\mu,\mu^{\prime}\in\mathcal{P}(\mathcal{X})$,
@@
Moreover, $\tilde{r}$ and $\tilde{g}$ are twice differentiable in $l\in{\mathbb{R}}^{d}$, and there exist $M_{1},M_{2}>0$ such that for all $x\in\mathcal{X},a\in\mathcal{A}$ and $l\in{\mathbb{R}}^{d}$,